OpenStudy (anonymous):

A bacteria culture starts with 1,500 bacteria and doubles in size every 4 hours. Find an exponential model for the size of the culture as a function of time t in hours

OpenStudy (anonymous):

a(t)=Ae^rt

OpenStudy (anonymous):

3000=1500e^4r

OpenStudy (anonymous):

ln2=4r

OpenStudy (anonymous):

t=ln2/4

OpenStudy (anonymous):

a(t)=Ae^(ln2)t/4

OpenStudy (anonymous):

sry... r = (ln2)/4 the last answer is still correct

OpenStudy (anonymous):

Valpey, what do you think?

OpenStudy (anonymous):

P = Noe^(kt) where P is amount after t time No is the initial amount . No is 1500 bacteria doubles in size every 4 hours so when t=4, P = 3000 3000 = 1500e^(4k) 2= e^(4k) ln(2) = 4k ln(2)/4 = k 0.17328679513999 =k so k=0.1733 P = 1500e^(0.1733t) so required equation is . F(t) = 1000e(0.34657t)

OpenStudy (anonymous):

initially its 1500 after 4hrs it becomes 1500*2 after 8 hrs----> 1500*2^2 after 12hrs---->1500*2^3 . . . after 4*t hrs---->1500*2^t so after t hrs its given by \[1500*2^t/4\]